OpenStudy (anonymous):

A flat rectangular piece of aluminum has a perimeter of 70 inches. The length is 11 inches longer than the width. Find the width.

OpenStudy (anonymous):

What is the formula for the perimeter of a rectangular?

OpenStudy (anonymous):

TO understand this question you need to know how to find the perimeter of a rectangle. Do you know the formula?

OpenStudy (anonymous):

w+L+w+L

OpenStudy (anonymous):

Yes or P=2L+2W

OpenStudy (anonymous):

The word problem gives us the permitter, which is P and it gives of the Length, which is 11 inches longer than the width. Length = (w + 11). So P = 2L + 2W P = 2(w + 11) + 2w can you finish and you see what I have done?

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

This will help you more P = 2(w + 11) + 2w 70 = 2(w + 11) + 2w Now finish 70 = 2(w + 11) + 2w

OpenStudy (anonymous):

23

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

I will work it out for you

OpenStudy (anonymous):

34

OpenStudy (anonymous):

70 = 2(w + 11) + 2w 70 = 2w + 22 + 2w 70 - 22 = 2w + 2w 48 = 4w 48/4 = 4w/4 12 = w w = 12

OpenStudy (anonymous):

I get it now thanks so much been a long day lol

OpenStudy (anonymous):

YW. Word problems can be tricky :-)